keil is going to mke 10 pounds of mixed nuts for a party. If Keil can spend $47.50 on nuts and peanuts cost $2.00 per pound and fancey nuts cost 7.50 per pound, how many pounds of each should he buy?
Answer by oberobic(2304)

p = lb of peanuts
200p = cost of peanuts in cents
f = lb of fancy nuts
750f = cost of fancy nuts in cents
.
4750 cents is the budget
.
200p + 750f = 4750
.
p + f = 10
so
p = 10-f
.
Substitute
.
200(10-f) + 750f = 4750
2000 -200f +750f = 4750
550f = 2750
f = 5
.
p = 10-f
p = 10-5
p = 5
.
Check the budget.
200p = 200*5 = 1000
750f = 750*5 = 3750
1000 + 3750 = 4750
Correct.
.
Answer: 5 lb of peanuts and 5 lb of fancy nuts will make 10 lb of mix at a cost of $47.50.
.
Done.